OpenStudy (anonymous):

what is the volume of a hemisphere with a radius of 10 ft

OpenStudy (anonymous):

the volume should be somewhere around 261.66 ft

OpenStudy (anonymous):

It's half of that of a sphere: \[V = \frac{4 \pi r^3}{3} / 2 = \frac{2 \pi r^3}{3}\] Insert r = 10 ft: \[V = \frac{2 \pi (10 \text{ ft})^3}{3} = \frac{2000\pi}{3} \text{ ft}^3\] We got different results, @TechnicianExpert.